Determine the frequency of the first overtone of a closed pipe of a length 30cm, if the end correction is 0.8cm and the velocity of sound is 346 meter per second
1
Expert's answer
2020-10-15T10:50:24-0400

End correction = 0.8 cm

The total assumed length = 30.8 cm

The frequency for the first overtone at a closed end pipe is

"f_3 = \frac{3V}{4L}"

"f_3 = \frac{3\times346\;m/s}{4\times 0.308\;m}= 842 \;Hz"

Answer: 842 Hz